OpenStudy (anonymous):

Two is thirty-one, divided by the sum of a number and four. Find the number. A. 11.5 B. –15.5 C. 13.5 D. none of the above

OpenStudy (studygurl14):

\[2 = \frac{31}{N + 4}\]

OpenStudy (anonymous):

THANKS

OpenStudy (studygurl14):

Solve for N

OpenStudy (anonymous):

2 = (31/(x+4)) 2*(x+4) = 31 x+4= (31/2) x = 15.5 - 4